You invested $400 million worth of PetroBras, the largest oil company in Brazil, in 2002. The stock price of PetroBras went from $5 in Brazilian Real in 2002 to $60 in Brazilian Real in 2008. At the same time, the Brasilian Real appreciated against the US by 200%. What is your return on investment in US Dollar term?
Answer #1

Assume that 1 USD = 10 BRL in 2002.

200% appreciation in real means that 1 USD = 10 / 3 = 3.33 BRL in 2008

Hence, total USD return per share = (USD Price in 2008) / (USD Price in 2002) - 1 = (60 / 3.33) / (5 / 10) - 1 = 3504%

Your $400 million investment would be worth $14,414 million.